a rectangle has sides (2x+3)cm and (4x+5 cm. how many squares of sides x can be cut from it?




Answer by stanbon(75887)

a rectangle has sides (2x+3)cm and (4x+5 cm. how many squares of sides x can be cut from it?
-------
Area = (2x+3)(4x+5) = 8x^2 +10x + 12x + 15 = 8x^2 + 22x + 15
-------
Ans: (8x^2+2x+15)/x^2 = 8 + Remainder
Ans: 8
==================
Cheers,
Stan H.